What is the mean number of convictions in 1990 and 2009?

Number of Public C 1990 2009 ot Level No of Convictions Govt Level No of Convictions Federal426 102 257 Federal 83 State 79 Stale Local 225 Local
Answer #1

Answer

We know that mean = (sum of all data values)/(total number of data values)

Calculation for 1990

we get, mean = (583+79+225)/3= 887/3 = 295.67

So, mean number of convictions in 1990 is 295.67

Calculation for 2009

mean = (sum of all data values)/(total number of data values)

we get, mean = (426+102+257)/3= 785/3 = 261.67

So, mean number of convictions in 2009 is 261.67